# Decoding Translation-Related Functional Sequences in 5 UTRs Using Interpretable Deep Learning Models

### v1

@@ Abstract
Understanding how 5’ untranslated regions (5’UTRs) regulate mRNA translation is critical for controlling protein expression and designing effective therapeutic mRNAs. While recent deep learning models have shown promise in predicting translational efficiency from 5’UTR sequences, most are constrained by fixed input lengths and limited interpretability. We introduce UTR-STCNet, a Transformer-based architecture for flexible and biologically grounded modeling of variable-length 5’UTRs. UTR-STCNet integrates a Saliency-Aware Token Clustering (SATC) module that iteratively aggregates nucleotide tokens into multi-scale, semantically meaningful units based on saliency scores. A Saliency-Guided Transformer (SGT) block then captures both local and distal regulatory dependencies using a lightweight attention mechanism. This combined architecture achieves efficient and interpretable modeling without input truncation or increased computational cost. Evaluated across three benchmark datasets, UTR-STCNet consistently outperforms state-of-the-art baselines in predicting mean ribosome load (MRL), a key proxy for translational efficiency. Moreover, the model recovers known functional elements such as upstream AUGs and Kozak motifs, highlighting its potential for mechanistic insight into translation regulation.
Introduction
@@
Messenger RNA (mRNA) therapeutics have emerged as a powerful platform for precise control of protein expression, with broad applications in vaccine development, cancer immunotherapy, and treatment of genetic disorders (Qin et al. 2022). A critical determinant of therapeutic efficacy is translation efficiency, which governs protein yield from synthetic mRNA constructs. The 5’ untranslated region (5’UTR) plays a central role in regulating translation initiation(Araujo et al. 2012), with sequence features such as upstream AUGs (uAUGs) (Wang and Rothnagel 2004) and Kozak (Li et al. 2017) motifs modulating ribosome scanning and start-site selection. Understanding the regulatory code embedded in 5’UTRs is thus essential for rational mRNA design (Cao et al. 2021).
@@
Recent advances in high-throughput experimental techniques, including ribosome profiling (Brar and Weissman 2015) and systematic mutagenesis (Kang et al. 2004), have produced large-scale datasets linking 5’UTR sequences to translational outcomes. These resources have enabled the application of deep learning to model translation regulation (Zhang et al. 2017), with mean ribosome load (MRL) frequently used as a key quantitative proxy for translation efficiency. Despite encouraging progress, current models face two critical limitations (Sample et al. 2019).
First, most architectures are trained on fixed-length inputs and lack the capacity to generalize to native 5’UTRs of varying lengths, requiring truncation that may obscure important regulatory elements. Second, existing models offer limited interpretability. Although some incorporate attention mechanisms or structural priors, they typically fail to attribute functional outcomes to specific sequence motifs such as uAUGs or Kozak elements (Kozak 1987), restricting their utility in mechanistic inference and variant prioritization.
@@
In addition, current models often struggle to scale in multi-task settings or to operate efficiently in resource-constrained environments, limiting their applicability in complex biological systems and real-world deployment scenarios.
@@
To address these challenges, we present UTR-STCNet, a Transformer-based deep learning framework for predicting translational efficiency from 5’UTR sequences. UTR-STCNet introduces two key innovations: a Saliency-Aware Token Clustering (SATC) module that groups and filters sequence tokens based on regulatory relevance, and a Saliency-Guided Transformer (SGT) block that models motif interactions across both local and long-range contexts using sparse attention. Our design enables accurate modeling of variable-length 5’UTRs with strong biological interpretability and computational efficiency. Our contributions are summarized as follows:
@@
  • We propose UTR-STCNet, a deep learning framework for predicting mean ribosome load from 5’UTR sequences, with strong generalization across datasets and biological contexts.
@@
Related Work
Self-Supervised Representation Learning for RNA Sequences
@@
A growing body of work has focused on learning contextual representations of RNA sequences through self-supervised pretraining, with the goal of capturing transcript-level regulatory information. Early methods relied on handcrafted k-mer features or shallow encoders, which lacked the capacity to model long-range dependencies and complex structural signals(Alipanahi et al. 2015). To address these limitations, RNA-specific language models such as RNA-BERT (Akiyama and Sakakibara 2022) and RNA-FM (Chen et al. 2022) were introduced, trained on large-scale transcriptomic datasets to produce contextualized RNA embeddings. These models significantly improved the modeling of splicing signals, secondary structure contexts, and untranslated region (UTR) motifs. In parallel, general-purpose contextual learning frameworks such as data2vec (Baevski et al. 2022) have been adapted to RNA domains, offering a unified modality-agnostic representation space. Building on this direction, mRNA2vec (Zhang et al. 2025) proposes a joint encoding strategy for 5’UTRs and coding regions by integrating both contextual and structural cues, enabling holistic modeling of mRNA molecules. Although genomic models like DNABERT have shown strong performance in DNA sequence tasks, their architecture and training focus limit their applicability to RNA-specific regulatory interpretation. Recent multimodal extensions incorporating structural or epigenomic data further expand the landscape of transferable RNA representations.
Supervised Deep Learning for 5’UTR-Based Translation Modeling
@@
In contrast to general-purpose representation learning, another line of work has focused on directly modeling translation efficiency from 5’UTR sequences using supervised deep learning. The 5’UTR plays a central role in post-transcriptional regulation, modulating translation initiation, ribosome recruitment, and start-site selection (Leppek, Das, and Barna 2018). Optimus (Sample et al. 2019) leverages global-pooling convolutional networks (CNNs) trained on mutagenesis data to predict mean ribosome load (MRL) from fixed-length sequences, capturing the functional impact of sequence variants. FramePool (Karollus, Avsec, and Gagneur 2021) extends this approach with a length-agnostic CNN architecture that enables frame-wise ribosome occupancy prediction, offering fine-resolution insights into initiation dynamics. To improve generalization across species and experimental conditions, MTrans (Zheng et al. 2023) applies a multi-task framework with shared CNN encoders and task-specific GRUs. UTR-LM (Chu et al. 2024) , in turn, pretrains a Transformer-based model on large-scale collections of 5’UTR sequences, learning transferable contextual embeddings for downstream translation-related tasks. Together, these models provide a strong foundation for 5’UTR-based translation modeling, demonstrating reliable performance across a range of predictive tasks. However, existing methods still face challenges in capturing interpretable, motif-aware dependencies and in modeling variable-length inputs without sacrificing biological relevance. These limitations motivate the development of specialized architectures that are both structurally adaptive and biologically transparent.
Method
@@
We present the architecture of UTR-STCNet in Figure.1. The model incorporates a basic Transformer block (Vaswani et al. 2017), a Saliency Awared Token Clustering (SATC) Module, a Saliency Guided Transformer (SGT) Block, and a multi-task prediction heads. SATC module hierarchically merges relevant tokens into saliency-aware representatives, enhancing structural compactness and semantic coherence. These refined tokens are subsequently processed by the SGT, which compresses the representations of 5’UTRs into more compact, saliency-weighted embeddings, thereby improving computational efficiency and reducing model complexity while preserving the ability to extract critical regulatory features and capture long-range dependencies. We adopt a multi-task learning framework where the feature extraction backbone is shared across all tasks. To capture task-specific characteristics, we design independent prediction heads for each task. Each task-specific head consists of a lightweight feedforward network that maps the shared representation to the task-specific output space. The task-specific heads then predict mean ribosome load based on the aggregated and interpretable representations.
Problem Formulation
@@
Figure 1: Overview of UTR-STCNet. The model takes variable-length 5’UTR sequences as input to predict mean ribosome load (MRL). It combines a Saliency-Aware Token Clustering (SATC) module for selecting informative tokens with a Saliency-Guided Transformer (SGT) block for modeling regulatory interactions. The output representations are biologically interpretable, enabling recovery of known sequence elements.
Saliency-Aware Token Clustering Module
Saliency Gating
$\alpha_{i}=\exp(\mathbf{w}^{\!\top}t_{i}).$ (1)
$D_{ij}=\frac{\|t_{i}-t_{j}\|_{2}}{\sigma\sqrt{d}},$ (2)
where $t_{i}\in\mathbb{R}^{d}$ denotes the $i$-th token embedding and $d$ is the feature dimension. The distance is normalized using the standard deviation of each embedding dimension, which reduces the dominance of high-variance features and improves the stability of distance computation. We then estimate the local density $\rho_{i}$ of each token using its $k$-nearest neighbors(Cover and Hart 1967):
where $\mathcal{N}_{k}(i)$ denotes the set of the $k$ nearest neighbors of token $i$. Then, for each token $i$, we explicitly compute a distance indicator $\delta_{i}$ representing the local density gap (Du, Ding, and Jia 2016) as:
$\quad\gamma_{i}=\rho_{i}\,\delta_{i},$ (5)
@@
Table 1: Performance Comparison of Single-Task Models Across Three 5’UTR Translation Datasets
@@
Saliency Guided Transformer Block
@@
$\begin{split}&\widetilde{\mathbf{K}}_{\ell}=\frac{\sum_{t\in\Gamma_{\ell}}p_{t}\odot k_{t}}{\sum_{t\in\Gamma_{\ell}}p_{t}+\varepsilon},\\ &\widetilde{\mathbf{V}}_{\ell}=\frac{\sum_{t\in\Gamma_{\ell}}p_{t}\odot v_{t}}{\sum_{t\in\Gamma_{\ell}}p_{t}+\varepsilon},\\ \end{split}$ (7)
@@
Although the aforementioned SATC module compresses the sequence along its sequence dimension, the feed-forward layer in standard transformers remains strictly position-wise and therefore struggles to capture short-range compositional patterns such as 3 to 5 nucleotide motifs, which are critical for 5’UTR regulation. To address this, we redesign the feed-forward sub-layer by incorporating a depth-wise convolutional structure (Chollet 2017) that effectively captures both known and novel motif signals, as well as other functionally relevant sequence elements. Specifically, the revised module applies a 1D depth-wise convolution followed by a GELU activation(Hendrycks and Gimpel 2016), with a residual connection to ensure gradient stability. Depth-wise convolutions are particularly well-suited for modeling local dependencies in sequence data, as they apply channel-specific filters to capture fine-grained patterns with high parameter efficiency. The projection weights before and after the convolution are shared across tokens, preserving computational efficiency while enabling each channel to extract distinct local motif features.
@@
This depth-wise convolutional mechanism introduces positional awareness at the motif level in a channel-efficient manner, complementing self-attention without requiring additional positional encodings. This mechanism further enables adaptation to varying input channel sizes following token aggregation, while preserving the spatial arrangement of tokens and capturing multi-scale features within the sequence. It integrates seamlessly with the channel-based token representations produced by the preceding SAG, enabling biologically meaningful context modeling into the SGT backbone. This enables the capture of functional differences across genes and complex regulatory contexts, enhances the modeling of dependencies between regulatory elements and sequence features, and ultimately improves the interpretability of functional signals in RNA sequences.
Experiments
Dataset
@@
We evaluated the proposed UTR-STCNet model using three publicly available 5’UTR translation datasets: MPRA-U, MPRA-H, and MPRA-V. These datasets were all derived from massively parallel reporter assays (MPRAs)(Cuperus et al. 2017), providing abundant sequence information and corresponding expression levels to support model training and validation. We first ranked all sequences based on total read counts obtained from sequencing, and selected the top 20,000 most highly expressed sequences from the MPRA-U dataset as test samples. Similarly, the top 7,600 highly expressed human-derived sequences from the MPRA-H dataset were selected as the test set. For the MPRA-V dataset, due to the variability in sequence lengths, we adopted a stratified random sampling strategy: sequences were divided into length bins ranging from 25 to 100 nucleotides, and from each bin, 100 randomly generated sequences and 30 natural human sequences were selected, resulting in a total of 9,880 test sequences. The remaining unselected sequences from each dataset were used to construct the combined training and validation sets.
Baselines
@@
  • Optimus (Sample et al. 2019) employs convolutional networks with residual blocks to forecast mean ribosome load across fixed-length 5’ UTRs.
  • FramePool (Karollus, Avsec, and Gagneur 2021) is a length-agnostic CNN baseline that predicts 5’UTR mean ribosome load using frame pooling.
  • RNA-FM(Chen et al. 2022) is a foundation model that captures RNA structure, 3D proximity, and RBP interactions from unannotated sequences.
  • MTtrans (Zheng et al. 2023) shares a convolutional encoder across tasks while task‑specific GRU towers yield length‑agnostic, cross‑assay translation‑rate predictions.
  • UTR-LM (Chu et al. 2024) is a transformer-based language model that enables unified functional modeling and pretraining over 5’UTR sequences.
  • mRNA2vec(Zhang et al., 2025) is a general-purpose pretrained model based on data2vec jointly encoding 5’UTR and CDS with contextual and structural signals for mRNA representation.
Performance Comparison

### v2

# Decoding Translation-Related Functional Sequences in 5’UTRs Using Interpretable Deep Learning Models

@@ Abstract
Understanding the regulatory role of 5’ untranslated regions (5’UTRs) in mRNA translation is critical for controlling protein expression and designing effective therapeutic mRNAs. While recent deep learning models have shown promise in predicting translational efficiency from 5’UTRs sequences, most are constrained by fixed input lengths and limited interpretability. We introduce UTR-STCNet, a Transformer-based architecture for flexible and biologically grounded modeling of variable-length 5’UTRs. UTR-STCNet integrates a Saliency-Aware Token Clustering (SATC) module that iteratively aggregates nucleotide tokens into multi-scale, semantically meaningful units based on saliency scores. A Saliency-Guided Transformer (SGT) block then captures both local and distal regulatory dependencies using a lightweight attention mechanism. This combined architecture achieves efficient and interpretable modeling without input truncation or increased computational cost. Evaluated across three benchmark datasets, UTR-STCNet consistently outperforms state-of-the-art baselines in predicting mean ribosome load (MRL), a key proxy for translational efficiency. Moreover, the model recovers known functional elements such as upstream AUGs and Kozak motifs, highlighting its potential for mechanistic insight into translation regulation.The code is available on GitHub [^1] https://github.com/Yu-Lab-Genomics/UTR-STCNet.
I Introduction
@@
Messenger RNA (mRNA) therapeutics have emerged as a powerful and promising platform for the precise, tunable, and designable control of protein expression, with broad applications in vaccine development, cancer immunotherapy, and the treatment of genetic disorders [21]. The ultimate efficacy of mRNA-based therapeutics is primarily determined by their translation efficiency, which dictates the amount and functionality of proteins synthesized from engineered mRNA constructs. Among the multiple factors influencing translation, the 5’ untranslated region (5’UTRs) serves as a key regulatory element that plays a central role in translation initiation [4]. Its sequence features, such as upstream AUGs (uAUGs) [25] and Kozak consensus motifs [20], can significantly affect ribosomal scanning dynamics and start-site selection, thereby modulating the overall efficiency of translation. Therefore, a systematic understanding of the regulatory code encoded within 5’UTRs is essential for the rational design and optimization of therapeutic mRNA constructs [7].
@@
Recent advances in high-throughput experimental techniques, including ribosome profiling [6] and systematic mutagenesis [15], have produced large-scale datasets that link 5’ untranslated region (5’UTRs) sequences to translational outcomes. These high-quality resources have enabled researchers to apply deep learning models to systematically explore the relationship between 5’UTRs sequences and translation efficiency [27], with mean ribosome load (MRL) frequently used as a key quantitative indicator of translation efficiency. The development of these models has not only deepened the quantitative understanding of translation regulation but also laid the groundwork for subsequent computational studies in mRNA design. Despite encouraging progress, current models still face two critical limitations [22].
First, most existing architectures are trained on fixed-length input sequences and therefore lack the ability to generalize to native 5’ untranslated regions (5’UTRs) of varying lengths. To accommodate such constraints, sequences are often truncated or padded to match the input window, which may mask or distort key regulatory elements. This preprocessing can lead to the loss of critical contextual information, especially in sequences containing long regulatory segments or multiple functional sites. As a result, models often struggle to capture length-dependent features present in natural 5’UTRs, reducing both prediction stability and generalization performance. Second, current models offer limited interpretability. Although some incorporate attention mechanisms or structural priors, they typically fail to assign functional effects to specific sequence motifs, such as upstream AUGs (uAUGs) or Kozak elements [18], thereby restricting their usefulness in mechanistic inference and variant prioritization. This limitation prevents models from providing biologically grounded explanations, leaving researchers reliant on overall prediction metrics rather than understanding which sequence features truly drive translation efficiency differences.
@@
In addition, current models often struggle to scale effectively in multi-task learning scenarios or to operate efficiently in resource-constrained environments, limiting their applicability in complex biological systems and real-world deployment contexts. In practical applications, such computational bottlenecks not only increase training and inference costs but also constrain the operational efficiency of models in high-throughput tasks.
@@
To address these challenges, we present UTR-STCNet, a Transformer-based deep learning framework for predicting translational efficiency from 5’ untranslated region (5’UTRs) sequences. UTR-STCNet introduces two key innovations: a Saliency-Aware Token Clustering (SATC) module that groups and filters sequence tokens based on regulatory relevance, and a Saliency-Guided Transformer (SGT) block that models motif interactions across both local and long-range contexts using sparse attention. This design enables the model to accurately capture translation-relevant signals from variable-length 5’UTRs while maintaining high biological interpretability and computational efficiency. Moreover, UTR-STCNet balances performance and scalability in its architectural design, exhibiting stable predictive performance across different data scales and task settings. The framework not only achieves significant improvements in prediction accuracy but also provides a new analytical perspective for understanding 5’UTRs regulatory mechanisms. Our contributions are summarized as follows:
@@
  • We propose UTR-STCNet, a deep learning framework for predicting mean ribosome load from 5’UTRs sequences, with strong generalization across datasets and biological contexts.
@@
II Related Work
II-A Self-Supervised Representation Learning for RNA Sequences
@@
A growing body of work has focused on learning contextual representations of RNA sequences through self-supervised pretraining, with the goal of capturing transcript-level regulatory information. Early methods relied on handcrafted k-mer features or shallow encoders, which lacked the capacity to model long-range dependencies and complex structural signals[2]. To address these limitations, RNA-specific language models such as RNA-BERT [1] and RNA-FM [8] were introduced, trained on large-scale transcriptomic datasets to produce contextualized RNA embeddings. These models significantly improved the modeling of splicing signals, secondary structure contexts, and untranslated region motifs. In parallel, general-purpose contextual learning frameworks such as data2vec [5] have been adapted to RNA domains, offering a unified modality-agnostic representation space. Building on this direction, mRNA2vec [26] proposes a joint encoding strategy for 5’UTRs and coding regions by integrating both contextual and structural cues, enabling holistic modeling of mRNA molecules. Although genomic models like DNABERT have shown strong performance in DNA sequence tasks, their architecture and training focus limit their applicability to RNA-specific regulatory interpretation. Recent multimodal extensions incorporating structural or epigenomic data further expand the landscape of transferable RNA representations.
II-B Supervised Deep Learning for 5’UTRs-Based Translation Modeling
@@
In contrast to general-purpose representation learning, another line of work has focused on directly modeling translation efficiency from 5’UTRs sequences using supervised deep learning. The 5’UTRs play a central role in post-transcriptional regulation, modulating translation initiation, ribosome recruitment, and start-site selection [19]. Optimus [22] leverages global-pooling convolutional networks (CNN) trained on mutagenesis data to predict MRL from fixed-length sequences, capturing the functional impact of sequence variants. FramePool [16] extends this approach with a length-agnostic CNN architecture that enables frame-wise ribosome occupancy prediction, offering fine-resolution insights into initiation dynamics. To improve generalization across species and experimental conditions, MTrans [28] applies a multi-task framework with shared CNN encoders and task-specific GRUs. UTR-LM [10] , in turn, pretrains a Transformer-based model on large-scale collections of 5’UTRs sequences, learning transferable contextual embeddings for downstream translation-related tasks. Together, these models provide a strong foundation for 5’UTRs-based translation modeling, demonstrating reliable performance across a range of predictive tasks. However, existing methods still face challenges in capturing interpretable, motif-aware dependencies and in modeling variable-length inputs without sacrificing biological relevance. These limitations motivate the development of specialized architectures that are both structurally adaptive and biologically transparent.
Figure 1: Overview of UTR-STCNet. The model takes variable-length 5’UTRs sequences as input to predict mean ribosome load (MRL). It combines a Saliency-Aware Token Clustering (SATC) module for selecting informative tokens with a Saliency-Guided Transformer (SGT) block for modeling regulatory interactions. The output representations are biologically interpretable, enabling recovery of known sequence elements.
III Method
@@
We present the architecture of UTR-STCNet in Figure.1. The model incorporates a basic Transformer block [24], a Saliency Awared Token Clustering (SATC) Module, a Saliency Guided Transformer (SGT) Block, and multi-task prediction heads. SATC module hierarchically merges relevant tokens into saliency-aware representatives, enhancing structural compactness and semantic coherence. These refined tokens are subsequently processed by the SGT, which compresses the representations of 5’UTRs into more compact, saliency-weighted embeddings, thereby improving computational efficiency and reducing model complexity while preserving the ability to extract critical regulatory features and capture long-range dependencies. We adopt a multi-task learning framework where the feature extraction backbone is shared across all tasks. To capture task-specific characteristics, we design independent prediction heads for each task. Each task-specific head consists of a lightweight feedforward network that maps the shared representation to the task-specific output space. The task-specific heads then predict mean ribosome load based on the aggregated and interpretable representations.
III-A Problem Formulation
@@
TABLE I: Performance Comparison of Single-Task Models Across Three 5’UTRs Translation Datasets
@@
III-B Saliency-Aware Token Clustering Module
Saliency Gating
$\alpha_{i}=\exp(\mathbf{w}^{\!\top}t_{i}).$ (1)
$D_{ij}=\frac{\|t_{i}-t_{j}\|_{2}}{\sigma\sqrt{d}},$ (2)
where $t_{i}\in\mathbb{R}^{d}$ denotes the $i$-th token embedding and $d$ is the feature dimension. The distance is normalized using the standard deviation of each embedding dimension, which reduces the dominance of high-variance features and improves the stability of distance computation. We then estimate the local density $\rho_{i}$ of each token using its $k$-nearest neighbors[11]:
where $\mathcal{N}_{k}(i)$ denotes the set of the $k$ nearest neighbors of token $i$. Then, for each token $i$, we explicitly compute a distance indicator $\delta_{i}$ representing the local density gap [13] as:
$\quad\gamma_{i}=\rho_{i}\,\delta_{i},$ (5)
@@ III-B Saliency-Aware Token Clustering Module
III-C Saliency Guided Transformer Block
@@
[…]
@@
Although the aforementioned SATC module compresses the sequence along its sequence dimension, the feed-forward layer in standard transformers remains strictly position-wise and therefore struggles to capture short-range compositional patterns such as 3 to 5 nucleotide motifs, which are critical for 5’UTRs regulation. To address this, we redesign the feed-forward sub-layer by incorporating a depth-wise convolutional structure [9] that effectively captures both known and novel motif signals, as well as other functionally relevant sequence elements. Specifically, the revised module applies a 1D depth-wise convolution followed by a GELU activation[14], with a residual connection to ensure gradient stability. Depth-wise convolutions are particularly well-suited for modeling local dependencies in sequence data, as they apply channel-specific filters to capture fine-grained patterns with high parameter efficiency. The projection weights before and after the convolution are shared across tokens, preserving computational efficiency while enabling each channel to extract distinct local motif features.
@@
This depth-wise convolutional mechanism introduces positional awareness at the motif level in a channel-efficient manner, complementing self-attention without requiring additional positional encodings. This mechanism further enables adaptation to varying input channel sizes following token aggregation, while preserving the spatial arrangement of tokens and capturing multi-scale features within the sequence. It integrates seamlessly with the channel-based token representations produced by the preceding SGA, enabling biologically meaningful context modeling into the SGT backbone. This enables the capture of functional differences across genes and complex regulatory contexts, enhances the modeling of dependencies between regulatory elements and sequence features, and ultimately improves the interpretability of functional signals in RNA sequences.
IV Experiments
IV-A Dataset
@@
We evaluated the proposed UTR-STCNet model using three publicly available 5’UTRs translation datasets: MPRA-U, MPRA-H, and MPRA-V. These datasets were all derived from massively parallel reporter assays (MPRAs)[12], providing abundant sequence information and corresponding expression levels to support model training and validation. We first ranked all sequences based on total read counts obtained from sequencing, and selected the top 20,000 most highly expressed sequences from the MPRA-U dataset as test samples. Similarly, the top 7,600 highly expressed human-derived sequences from the MPRA-H dataset were selected as the test set. For the MPRA-V dataset, due to the variability in sequence lengths, we adopted a stratified random sampling strategy: sequences were divided into length bins ranging from 25 to 100 nucleotides, and from each bin, 100 randomly generated sequences and 30 natural human sequences were selected, resulting in a total of 9,880 test sequences. The remaining unselected sequences from each dataset were used to construct the combined training and validation sets.
IV-B Baselines
@@
  • Optimus [22] employs convolutional networks with residual blocks to forecast mean ribosome load across fixed-length 5’ UTRs.
  • FramePool [16] is a length-agnostic CNN baseline that predicts 5’UTRs mean ribosome load using frame pooling.
  • RNA-FM[8] is a foundation model that captures RNA structure, 3D proximity, and RBP interactions from unannotated sequences.
  • MTtrans [28] shares a convolutional encoder across tasks while task‑specific GRU towers yield length‑agnostic, cross‑assay translation‑rate predictions.
  • UTR-LM [10] is a transformer-based language model that enables unified functional modeling and pretraining over 5’UTRs sequences.
  • mRNA2vec(Zhang et al., 2025) is a general-purpose pretrained model based on data2vec jointly encoding 5’UTRs and CDS with contextual and structural signals for mRNA representation.
IV-C Performance Comparison